# Data-driven $h^{2}$ model reduction for linear discrete-time systems

### v2

@@ Abstract
We present a new framework of $h^{2}$ optimal model reduction for linear discrete-time systems. Our main contribution is to create optimal reduced-order models in the $h^{2}$-norm sense directly from the measurement data alone, without using the information of the original system. In particular, we focus on the fact that the gradient of the $h^{2}$ model reduction problem is expressed using the discrete-time Lyapunov equation and the discrete-time Sylvester equation, and derive the data-driven gradient. In the proposed algorithm, the initial point can be chosen as the output of the existing data-driven methods. In numerical experiments, we demonstrate that, for a modeling task in neuroscience, our method constructs a reduced-order model that outperforms DMDc in terms of the $h^{2}$-norm.
@@ I Introduction
Data-driven model reduction constructs the ROMs that are capable of reproducing the input-output behavior of large-scale dynamical systems with high accuracy directly from the measurement data alone. Unlike traditional approaches, an explicit parametric system-identification step is not required; therefore, no additional identification-model error is introduced. Several well-known data-driven MOR methods have been proposed due to the increasing use of data. The data-driven balanced truncation method [13, 14, 15, 16] is one of the common data-driven model reduction inspired by the balanced truncation method, which is a model-based MOR. In this method, balanced truncation is performed by using the measurement data. For example, [15] performs MOR by estimating the Gramians from the data information. Another well-known approach is data-driven interpolatory model reduction. This approach includes the Loewner framework [17, 18, 19, 20, 21, 22, 23] and the AAA algorithm [24, 25]. Other data-driven methods, such as proper orthogonal decomposition (POD) [26], dynamic mode decomposition (DMD) [27, 28], or operator inference [29, 30] can also be employed. These approaches are the leading data-driven model reduction, but they may not yield optimal ROMs in the $h^{2}$-norm sense. In such cases, the output of the resulting reduced system may not approximate that of the original system.
In this paper, we propose a data-driven model reduction method for linear discrete-time systems. We focus on obtaining the optimal ROMs in the $h^{2}$-norm sense, using the measurement time-domain data. The contributions of this research are the following two points:
1) We propose a new framework for data-driven model reduction for linear discrete-time systems. We focus on obtaining the ROMs that perform well in terms of the $h^{2}$-norm, directly from the measurement data. In our approach, there is no need to use system identification methods, as in traditional approaches. Furthermore, under certain assumptions about the measurement data, we derive the gradients of the optimization problem characterized by the data.
2) We develop an algorithm for the proposed method and conduct numerical experiments. The proposed algorithm can be hybridized with existing data-driven MOR. In numerical experiments, we demonstrate that, for a modeling task in neuroscience, our method constructs a ROM that outperforms DMDc in terms of the $h^{2}$-norm.
@@
Notation: We denote the imaginary unit by $\mathrm{i}$. We also denote the Frobenius norm, transpose, trace, the Moore-Penrose pseudo-inverse, the $i$-th eigenvalue, the $i$-th row vector of the matrix $A$, and the $(i,j)$-th element of the matrix $A$ by $\|A\|_{F}$, $A^{\top}$, $\operatorname{tr}A$, $A^{{\dagger}}$, $\lambda_{i}(A)$, and $A(i,:)$, $[A]_{i,j}$, respectively. In addition to that, we denote the complex conjugate transpose of $A\in\mathbb{C}^{n\times n}$ by $A^{*}\in\mathbb{C}^{n\times n}$.
@@ II Preliminaries
In this section, we summarize a model-based $h^{2}$ MOR [4, 7, 9, 10] for linear time-invariant (LTI) discrete-time dynamical systems
@@
with transfer function $H(z):=C(zI_{n}-A)^{-1}B$, where $k$ indicates the temporal iteration and $u_{k}\in\mathbb{R}^{m}$, $x_{k}\in\mathbb{R}^{n}$, and $y_{k}\in\mathbb{R}^{p}$ denote the input, state, and output at time $t_{k}$, respectively. The matrices $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, and $C\in\mathbb{R}^{p\times n}$ are constant matrices, which are known.
In this study, we assume that $C=I_{n}$. This means $p=n$. That is, we consider a system in which all state data are observed as outputs. This assumption is commonly made in the context of DMD, as explained in Section III. In addition, we assume that the system $\Sigma$ is asymptotically stable, i.e. all eigenvalues of $A\in\mathbb{R}^{n\times n}$ lie inside the unit circle. Such a matrix $A$ is called a stable matrix.
The ROM of system (1) is defined as
@@
with transfer function $\hat{H}(z):=\hat{C}(zI_{r}-\hat{A})^{-1}\hat{B}$, where $\hat{x}_{k}\in\mathbb{R}^{r}$, $\hat{y}_{k}\in\mathbb{R}^{n}$, and $r\ll n$. In this context, $\hat{C}$ can be regarded as a matrix that reconstructs the dimensionally reduced state data back into the original output space, and it is employed to compare original system (1) and ROM (2).
II-A Model-Based $h^{2}$ Model Order Reduction
To track the output error $x_{k}-\hat{C}\hat{x}_{k}$ between (1) and (2), we focus on finding a ROM (2) that minimizes the approximation error $H-\hat{H}$ in the $h^{2}$-norm under the constraints that $(\hat{A},\hat{B},\hat{C})\in\mathbb{R}^{r\times r}\times\mathbb{R}^{r\times m}\times\mathbb{R}^{n\times r}$ and $\hat{A}$ is stable. Note that the $h^{2}$-norm of the discrete-time system $\Sigma$ is defined as
@@ II-A Model-Based $h^{2}$ Model Order Reduction
The $h^{2}$ MOR problem for the original system (1) is written as
$&\mathop{\rm minimize}\limits&&\|H-\hat{H}\|_{h^{2}}\\ &\mathop{\rm subject\ to}\limits&&(\hat{A},\hat{B},\hat{C})\in\mathbb{R}^{r\times r}\times\mathbb{R}^{r\times m}\times\mathbb{R}^{n\times r},\\ &&&\hat{A}\>\text{is stable}.\end{aligned}$ (3)
Proposition 1 ([4][10]).
Consider the asymptotically stable systems (1) with $C=I_{n}$ and (2). Then, the objective function of (3) can be rewritten as
$\|H-\hat{H}\|_{h^{2}}^{2}$ $=\operatorname{tr}(\Sigma_{c})+f(\hat{A},\hat{B},\hat{C})$
$=\operatorname{tr}(B^{\top}\Sigma_{o}B)+f(\hat{A},\hat{B},\hat{C}).$
Here,
$f(\hat{A},\hat{B},\hat{C})$ $:=\operatorname{tr}(\hat{C}P\hat{C}^{\top})-2\operatorname{tr}(R\hat{C}^{\top})$
$=\operatorname{tr}(\hat{B}^{\top}Q\hat{B})+2\operatorname{tr}(B^{\top}S\hat{B}),$
@@
From Proposition 1, (3) can be rewritten as
$&\mathop{\rm minimize}\limits&&f(\hat{A},\hat{B},\hat{C})\\ &\mathop{\rm subject\ to}\limits&&(\hat{A},\hat{B},\hat{C})\in\mathbb{R}^{r\times r}\times\mathbb{R}^{r\times m}\times\mathbb{R}^{n\times r},\\ &&&\hat{A}\>\text{is stable}.\end{aligned}$ (10)
A modification as in (10) eliminates the need to compute the large-Lyapunov equations as in (4) and (5).
@@ II-B Gradients of (10)
The gradients of a real scalar smooth function $f(X)$ of a real matrix variable $X\in\mathbb{R}^{n\times p}$ is the real matrix $\nabla_{X}f(X)\in\mathbb{R}^{n\times p}$ defined by
$[\nabla_{X}f(X)]_{i,j}=\frac{\partial}{\partial X_{i,j}}f(X),i=1,\ldots,n,j=1,\ldots,p.$
@@
When the gradients $(\nabla_{\hat{A}}f,\nabla_{\hat{B}}f,\nabla_{\hat{C}}f)$ are all zero, the stationary conditions derived in [10], i.e., Wilson’s first order necessary conditions, are satisfied. In this study, we derive these gradients directly from the measurement data without using the constant matrices $A$ and $B$ of the system (1).
@@ III Problem formulation
We formulate the problem addressed in this study under the assumption that the matrices $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ for the original system (1) are unknown. Note that this is a different setting from Section II. Instead, assume that $N$ measurement data sets obtained from the true system (1) is given by
$\begin{pmatrix}x_{1,1}&x_{1,2}&\dots&x_{1,L}\\ x_{2,1}&x_{2,2}&\dots&x_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ x_{N,1}&x_{N,2}&\dots&x_{N,L}\end{pmatrix},\quad\begin{pmatrix}u_{1,1}&u_{1,2}&\dots&u_{1,L}\\ u_{2,1}&u_{2,2}&\dots&u_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ u_{N,1}&u_{N,2}&\dots&u_{N,L}\end{pmatrix}$ (14)
@@
Problem 1 is important for the following reason: we examine a system with transfer function $H_{\textrm{SI}}$, which is obtained by a system identification method. Then,
$\|H-\hat{H}\|_{h^{2}}$ $=\|H-H_{\textrm{SI}}+H_{\textrm{SI}}-\hat{H}\|_{h^{2}}$
$\leq\|H-H_{\textrm{SI}}\|_{h^{2}}+\|H_{\textrm{SI}}-\hat{H}\|_{h^{2}}.$ (16)
In traditional approaches, the minimization of the error $\|H_{\textrm{SI}}-\hat{H}\|_{h^{2}}$ is considered. However, if errors $\|H-H_{\textrm{SI}}\|_{h^{2}}$ arise from system identification, minimizing $\|H_{\textrm{SI}}-\hat{H}\|_{h^{2}}$ may still not result in satisfactory ROMs in the $h^{2}$-norm. Furthermore, if the system identified is unstable, $\|H_{\textrm{SI}}-\hat{H}\|_{h^{2}}$ cannot be defined and the $h^{2}$ MOR may not be performed. To overcome these problems, this paper proposes a MOR method to optimize the left-hand side of (III) directly from the measurement data.
@@ IV Data-driven $h^{2}$ Model Reduction for Discrete-time Systems
In this section, we describe a data-driven $h^{2}$ MOR for (10). By making some assumptions on the measurement data of the original system (1), we show that ROMs can be obtained directly from the data when the matrices $A$ and $B$ of (1) are unknown.
@@ IV-A Data-driven Reconstruction of the State Equation
$X_{\mathrm{p}}:=$ $\big{(}x_{1,2},x_{1,3},\ldots,x_{1,L},x_{2,2},x_{2,3},\ldots,x_{2,L},$
@@
where the $i$-th row of $X_{\mathrm{p}}$ corresponds to the $i$-th row of $X$, shifted by one step according to the system’s state equation.
Next, we represent the system’s state equation using the measurement data. Using (15) and the data matrices (17),
@@
Note that $Z_{A}$ corresponds to the matrix of state data for the dual system in (15).
In general, the solutions $Z_{A}$, $Z_{B}$, and $U_{B}$ of (20) and (23) are not uniquely determined as in (21), (22), and (24). We denote the solutions obtained by solving (20) and (23) as $Z_{A,\text{data}}$, $Z_{B,\text{data}}$, and $U_{B,\text{data}}$, respectively.
To prove the uniqueness, we consider the following assumption.
@@
    The eigenvalues of $\hat{A}$ are inside the unit circle and strictly positive, i.e, $\forall i\in\{1,2,\ldots,r\},0<|\lambda_{i}(\hat{A})|<1$.
@@
    $\operatorname{rank}\begin{bmatrix}X&U\end{bmatrix}=n+m$.
@@
Remark 2.
Assumptions (A1)-(A6) may be easily achieved for the following reasons.
Assumptions (A1)-(A3) are conditions related to the eigenvalues of the $\hat{A}$ matrix. As described in Section V, the proposed algorithm allows the $\hat{A}$ matrix to be constructed to satisfy these assumptions. For details, refer to Section V.
Assumptions (A4)-(A6) are specifically related to the measurement data. If the number of data sets $N$ and measurement instances $L$ satisfy $NL-N\geq n+m$, and the vectors $(x_{i,k},u_{i,k})^{\top}$, $x_{i,k}$, and $u_{i,k}$ are linearly independent for each $i=1,\ldots,N$ and $k=1,\ldots,L$, respectively, then the assumptions hold. Furthermore, these assumptions are related to the persistency of excitation condition of the input [35]. Specifically, when the input $u$ possesses sufficient persistent excitation, the Hankel matrix of $U$ becomes full rank, and the rank of $X$ coincides with the state dimension $n$. Therefore, for example, as shown in [36], if an input $u$ that satisfies the persistency of excitation condition is generated, both $X$ and $U$ are guaranteed to be full rank.

### v3

@@ Abstract
We present a data-driven framework for $h^{2}$-optimal model reduction for linear discrete-time systems. Our main contribution is to create optimal reduced-order models in the $h^{2}$-norm sense directly from the measurement data alone, without using the information about the original system. In particular, we focus on the fact that the gradients of the $h^{2}$ model reduction problem are expressed using the discrete-time Lyapunov equation and the discrete-time Sylvester equation, and derive the data-driven gradients. The proposed algorithm uses the output of an existing MOR as the initial point, and convergence to a stationary point is guaranteed under certain assumptions. In numerical experiments, we demonstrate that, for a modeling task in neuroscience, our method constructs a reduced-order model that outperforms DMDc in terms of the $h^{2}$-norm.
@@ I Introduction
Data-driven model reduction constructs the ROMs that are capable of reproducing the input-output behavior of large-scale dynamical systems with high accuracy directly from the measurement data alone. Several well-known data-driven MOR methods have been proposed due to the increasing use of data. The data-driven balanced truncation (BT) [13, 14, 15, 16] is one of the common data-driven model reduction inspired by the BT, which is a model-based MOR. In this method, BT is performed by using the measurement data. For example, [15] performs MOR by estimating the Gramians from the data information. Another well-known approach is data-driven interpolatory model reduction. This approach includes the Loewner framework  [17, 18] and the AAA algorithm [19, 20]. Other data-driven methods, such as dynamic mode decomposition (DMD) [21, 22] or operator inference [23, 24] can also be employed. These approaches are the leading data-driven model reduction, but they may not yield optimal ROMs in the $h^{2}$-norm sense. In such cases, the output of the resulting reduced system may not approximate that of the original system.
In this paper, we propose a data-driven model reduction method for linear discrete-time systems. We focus on obtaining the optimal ROMs in the $h^{2}$-norm sense, using the measurement time-domain data. The contributions of this research are:
1) We propose a new framework for data-driven model reduction for linear discrete-time systems. We focus on obtaining the ROMs that perform well in terms of the $h^{2}$-norm, directly from the measurement data. Furthermore, under certain assumptions about the measurement data, we derive the gradients of the optimization problem characterized by the data.
2) We develop an optimization algorithm for the proposed method. The proposed algorithm can be hybridized with existing MOR. Furthermore, convergence to a stationary point is guaranteed under certain assumptions. We demonstrate that, for a modeling task in neuroscience, our method constructs a ROM that outperforms DMDc in terms of the $h^{2}$-norm.
@@
Notation: We denote the imaginary unit by $\mathrm{i}$. We also denote the 2-norm of a vector $a$, and for a matrix $A$ its Frobenius norm, transpose, complex conjugate transpose (for complex $A$), trace, Moore–Penrose pseudoinverse, $i$th eigenvalue, $i$th row, and $(i,j)$th entry by $\|a\|$, $\|A\|_{F}$, $^{\top}, AA^{*}$, $\operatorname{tr}A$, $A^{\dagger}$, $\lambda_{i}(A)$, $A(i,:)$, and $A]_{i,j}$, respectively.
@@ II Preliminaries
In this section, we summarize a model-based $h^{2}$ MOR [4, 7, 9, 10] for linear time-invariant discrete-time systems
$\begin{aligned} x_{k+1}&=Ax_{k}+Bu_{k},\\ y_{k}&=Cx_{k}\end{aligned}$ (1)
with transfer function $H(z):=C(zI_{n}-A)^{-1}B$, where $z\in\mathbb{C}$ is the $z$-transform variable, $k$ is the time index, and $u_{k}\in\mathbb{R}^{m}$, $x_{k}\in\mathbb{R}^{n}$, and $y_{k}\in\mathbb{R}^{p}$ denote the input, state, and output at time $t_{k}$, respectively. The matrices $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, and $C\in\mathbb{R}^{p\times n}$ are constant matrices, which are known. Furthermore, the $h^{2}$-norm of the system (1) is defined as
$\|H\|_{h^{2}}:=\left(\frac{1}{2\pi}\int_{0}^{2\pi}\operatorname{tr}\left[H^{*}(e^{-\mathrm{i}\theta})H(e^{\mathrm{i}\theta})\right]d\theta\right)^{\frac{1}{2}}.$
In this study, we assume that $C=I_{n}$. This means $p=n$. That is, we consider a system in which all state data are observed as outputs. This assumption is commonly made in the context of DMD, as explained in Section III. In addition, we assume that the system (1) is asymptotically stable, i.e. all eigenvalues of $A\in\mathbb{R}^{n\times n}$ lie inside the unit circle. Such a matrix $A$ is called a stable matrix.
The ROM of the system (1) is defined as
$\begin{aligned} \hat{x}_{k+1}&=\hat{A}\hat{x}_{k}+\hat{B}u_{k},\\ \hat{y}_{k}&=\hat{C}\hat{x}_{k}\end{aligned}$ (2)
with transfer function $\hat{H}(z):=\hat{C}(zI_{r}-\hat{A})^{-1}\hat{B}$, where $\hat{x}_{k}\in\mathbb{R}^{r}$, $\hat{y}_{k}\in\mathbb{R}^{n}$, and $r\ll n$. Following [7, 12], we obtain
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\max_{k\in\mathbb{Z}}\|x_{k}-\hat{C}\hat{x}_{k}\|\leq\|H-\hat{H}\|_{h^{2}}\cdot\sqrt{\sum_{k\in\mathbb{Z}}\|u_{k}\|^{2}}.$} (3)
(3) means that, under sufficiently small $\|u_{k}\|$, the output error $\|x_{k}-\hat{C}\hat{x}_{k}\|$ becomes smaller by minimizing the approximation error $\|H-\hat{H}\|_{h^{2}}$.
In this context, $\hat{C}$ can be regarded as a matrix that reconstructs the dimensionally reduced state data back into the original output space, and it is employed to compare the original system (1) and ROM (2).
II-A Model-Based $h^{2}$ Model Order Reduction
Motivated by (3), we focus on finding a ROM (2) that minimizes $\|H-\hat{H}\|_{h^{2}}$ so as to bound the output error $\|x_{k}-\hat{C}\hat{x}_{k}\|$, under the constraint $(\hat{A},\hat{B},\hat{C})\in\mathcal{E}_{s}$, where
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{E}_{s}:=\{(\hat{A},\hat{B},\hat{C})\in\mathbb{R}^{r\times r}\times\mathbb{R}^{r\times m}\times\mathbb{R}^{n\times r}\mid\hat{A}\ \text{is stable}\}.}$ (4)
Problem 1.
The $h^{2}$ MOR problem for (1) is
$\mathop{\rm minimize}\limits$ $\|H-\hat{H}\|_{h^{2}}$
$\mathop{\rm subject\ to}\limits$ $(\hat{A},\hat{B},\hat{C})\in\mathcal{E}_{s}.$
Proposition 1 ([4][10]).
Consider the asymptotically stable systems (1) with $C=I_{n}$ and (2). Then, the objective function of Problem 1 can be rewritten as
@@ II-A Model-Based $h^{2}$ Model Order Reduction
Here,
@@
[…]
@@
From Proposition 1, Problem 1 can be rewritten as
$\begin{aligned} &\mathop{\rm minimize}\limits&&f(\hat{A},\hat{B},\hat{C})\\ &\mathop{\rm subject\ to}\limits&&(\hat{A},\hat{B},\hat{C})\in\mathcal{E}_{s}.\end{aligned}$ (11)
A modification as in (11) eliminates the need to compute the large Lyapunov equations as in (5) and (6).
@@ II-B Gradients of (11)
The gradient of a real scalar smooth function $f(X)$ of a real matrix variable $X\in\mathbb{R}^{n\times p}$ is the real matrix $\nabla_{X}f(X)\in\mathbb{R}^{n\times p}$ defined by
$[\nabla_{X}f(X)]_{i,j}=\frac{\partial}{\partial X_{i,j}}f(X),\>\>i=1,\ldots,n,\>\>j=1,\ldots,p.$
@@
Since the set of stable matrices is open [25], a stationary point can be defined as follows.
Definition 2.
[…]
When the gradients $(\nabla_{\hat{A}}f,\nabla_{\hat{B}}f,\nabla_{\hat{C}}f)$ are all zero at certain points, the stationary conditions derived in [10], i.e., Wilson’s first order necessary conditions, are satisfied. However, note that these points may not necessarily belong to the feasible set ${\mathcal{E}}_{\mathrm{s}}$. In this study, we derive these gradients directly from the measurement data without using the constant matrices $A$ and $B$ of the system (1), and propose a gradient-based optimization algorithm for (11).
@@ III Problem formulation
We formulate the problem addressed in this study under the assumption that the matrices $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ for the original system (1) are unknown. Note that this is a different setting from Section II. Instead, assume that $N$ measurement data sets obtained from the true system (1) are given by
$\begin{pmatrix}x_{1,1}&x_{1,2}&\dots&x_{1,L}\\ x_{2,1}&x_{2,2}&\dots&x_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ x_{N,1}&x_{N,2}&\dots&x_{N,L}\end{pmatrix},\>\>\begin{pmatrix}u_{1,1}&u_{1,2}&\dots&u_{1,L}\\ u_{2,1}&u_{2,2}&\dots&u_{2,L}\\ \vdots&\vdots&\ddots&\vdots\\ u_{N,1}&u_{N,2}&\dots&u_{N,L}\end{pmatrix}$ (15)
@@
Problem 2 is important for the following reason: we examine a system with transfer function $H_{\textrm{SI}}$, which is obtained by system identification. By the triangle inequality,
$\|H-\hat{H}\|_{h^{2}}$ $\leq\|H-H_{\textrm{SI}}\|_{h^{2}}+\|H_{\textrm{SI}}-\hat{H}\|_{h^{2}}.$ (17)
When using classical system identification [30, 31], the error $\|H_{\textrm{SI}}-\hat{H}\|_{h^{2}}$ is taken into account. However, if errors $\|H-H_{\textrm{SI}}\|_{h^{2}}$ arise from system identification, minimizing $\|H_{\textrm{SI}}-\hat{H}\|_{h^{2}}$ may still not result in satisfactory ROMs in the $h^{2}$-norm. Furthermore, if the system identified is unstable, $\|H_{\textrm{SI}}-\hat{H}\|_{h^{2}}$ cannot be defined and the $h^{2}$ MOR may not be performed. To overcome these problems, this paper proposes a MOR method to optimize the left-hand side of (17) directly from the measurement data.
@@ IV Data-driven $h^{2}$ Model Reduction for Discrete-time Systems
We describe a data-driven $h^{2}$ MOR for (11). By making some assumptions on the measurement data of the original system (1), we show that ROMs can be obtained directly from the data when the matrices $A$ and $B$ of (1) are unknown.
@@ IV-A Data-driven Reconstruction of the State Equation
$X_{\mathrm{p}}$= $\big(x_{1,2},x_{1,3},\ldots,x_{1,L},x_{2,2},x_{2,3},\ldots,x_{2,L},$
@@
where the $i$th row of $X_{\mathrm{p}}$ corresponds to the $i$th row of $X$, shifted by one step according to the system’s state equation.
Using (16) and the data matrices (18),
@@
Note that $Z_{A}$ corresponds to the matrix of state data for the dual system in (16). In general, the solutions $Z_{A}$, $Z_{B}$, and $U_{B}$ of (21) and (23) are not uniquely determined as in (22) and (24). We denote the solutions obtained by solving (21) and (23) as $Z_{A,\text{data}}$, $Z_{B,\text{data}}$, and $U_{B,\text{data}}$, respectively. To prove uniqueness, we assume the following.
@@
    The eigenvalues of $\hat{A}$ are inside the unit circle and strictly positive, i.e., $\forall i\in\{1,2,\ldots,r\},0<|\lambda_{i}(\hat{A})|<1$.
@@
    $\operatorname{rank}\!\begin{pmatrix}X&U\end{pmatrix}=n+m$.
@@
(A1)-(A6) may be easily achieved for the following reasons. (A1)-(A3) are conditions related to the eigenvalues of $\hat{A}$. As described in Section V, the proposed algorithm allows $\hat{A}$ to be constructed to satisfy these assumptions. For details, refer to Section V.
(A4)-(A6) are specifically related to the measurement data. If the number of data sets $N$ and measurement instances $L$ satisfy $NL-N\geq n+m$, and the vectors $(x_{i,k},u_{i,k})^{\top}$, $x_{i,k}$, and $u_{i,k}$ are linearly independent for each $i=1,\ldots,N$ and $k=1,\ldots,L$, respectively, then the assumptions hold. Furthermore, these assumptions are related to the persistency of excitation condition of the input [32]. Specifically, when the input $u$ possesses sufficient persistent excitation, the Hankel matrix of $U$ becomes full rank, and the rank of $X$ coincides with the state dimension $n$. Therefore, for example, as shown in [33], if an input $u$ that satisfies the persistency of excitation condition is generated, both $X$ and $U$ are guaranteed to be full rank.